please help me solve these problems, Thank You!:
(2a^2n^5)(-3an^2)=
(3qn^2)^3=
(y^7/6)^2=

Answer by checkley77(12569)

2A^2N^5
-3AN^2 MULTIPLY.
------------------
6A^3N^7 ANSWER.
************************************
(3QN^2)^3=27Q^3N^6 ANSWER
************************************
(Y^7/6)^2=Y^14/6=Y^7/3 ANSWER.